OpenStudy (anonymous):

how would I solve and simplify -(-7y+12)

OpenStudy (anonymous):

7y-12=0 7y=12 y=12/7

OpenStudy (anonymous):

are you sure ?

OpenStudy (skullpatrol):

No, it is an expression, NOT an equation

OpenStudy (anonymous):

so it would be -5

OpenStudy (skullpatrol):

$$\Huge -(-7y+12)=-1 (-7y+12)$$

OpenStudy (skullpatrol):

now use the distributive property to simplify

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (anonymous):

I don't know what that means

OpenStudy (anonymous):

im bad at math its my worst subject

OpenStudy (skullpatrol):